OpenStudy (anonymous):

How many different sequences of 4 digits are possible if the first digit must be 3, 4, or 5 and if the sequence may not end in 000? Repetition of digits is allowed. A. 1,512 B. 2,999 C. 2,997 D. 2,000

OpenStudy (anonymous):

my guess is \(3\times 10\times 10\times 10-1\)

OpenStudy (anonymous):

3,000 - 3=2997